# Many-body localization crossover is sharper in quasiperiodic spin chains

### v1

@@ Abstract
Due to a phenomenon of many-body localization (MBL), the strong disorder may significantly slow down or even completely hinder the thermalization of quantum many-body systems. A sufficiently deep quasiperiodic potential may also inhibit thermalization. In this work, we numerically demonstrate direct differences in the behavior of standard ergodicity-breaking indicators at the MBL crossover in random and quasiperiodic systems. Our key finding is the exponential increase in the sharpness of the MBL crossover with system size for quasiperiodic systems, a trend that is only linear in disordered systems. The strong tendency towards a non-analytic behavior in quasiperiodic systems is consistent with the existence of dynamical regimes with sharply defined boundaries or an MBL phase transition. It highlights the importance of quasiperiodic systems for our understanding of many-body dynamics.
Introduction.
@@
where $\kappa$ is an incommensurability factor (e.g. $\kappa=\kappa_{\mathrm{gr}}\equiv(\sqrt{5}-1)/2$) and $\varphi$ is a global phase randomly selected from the interval $[0,2\pi]$. Increase in the amplitude $W$ of the RD and QP potential leads to inhibition of thermalization and the onset of MBL regime [28, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92]. The phenomenology of MBL regime in QP systems is very similar to the features of MBL in the presence of RD [93, 94, 95, 96]. The former are realized naturally in optical lattices by imposing a secondary standing wave with a period incommensurate with the primary lattice [97, 98, 99, 100, 101]. However, the two types of potentials are starkly different. The fields $h_{i}$ for RD are uncorrelated, while the QP potential (2) is deterministically fixed by the single phase $\varphi$, resulting in non-vanishing correlations between the on-site potential values at arbitrarily distant sites. Consequently, RD systems are more heterogeneous and host Griffiths type [102] rare regions of locally weaker disorder [103, 104, 105, 106, 107, 108, 109, 110, 111], while QP systems are more uniform [112].
@@
Do the disparities of the RD and QP potential translate to any essential differences in the properties of MBL in the two types of systems? An affirmative answer is suggested by quantities measuring sample-to-sample fluctuations of entanglement entropy [113], level statistics [114], and properties of states in the basis of Hilbert space [115], which indicate smaller fluctuations of system’s properties for QP potential than for RD. Similarly, finite-size collapses point out weaker system size drifts for QP potential [116], simulations of time dynamics reveal weak but noticeable long-time oscillations in correlation functions of local observables for QP potential, unobservable for the RD case [52], and diffusion constant vanishes faster in QP chains [117]. Nevertheless, direct numerical evidence for standard ergodicity-breaking indicators demonstrating the difference between MBL in RD and QP potential is missing, and the role of the statistical disparities between the two potentials for the dynamics may be questioned [118].
@@
In this work, we perform a large-scale numerical investigation of ETH-MBL crossover in QP and RD Floquet systems and support further our conclusions with results for RD (1) and QP (2) autonomous spin chains. Analyzing QP models, we find persistent drifts of ergodicity-breaking indicators at the ETH-MBL crossover, akin to the RD case. Despite these similarities, there is an immediate disparity between the QP and RD potentials. The ETH-MBL crossover sharpens up exponentially quickly with increasing system size $L$ for QP potential, while the sharpening up is only linear with $L$ for the RD. This demonstrates a direct difference between the standard bulk ergodicity-breaking indicators at ETH-MBL crossover in the RD and QP systems.
Models.
@@
Similarly to the autonomous spin chains, periodically driven Floquet systems may underro MBL [119, 120, 121, 122, 123, 124, 125, 126, 127] which prevents their heating [128], enabling creation of driven time crystals [129, 130, 131, 132, 133, 134, 135] and stabilization of Floquet insulators [136, 137, 138, 139, 140]. We focus on a Kicked Ising Model (KIM[141, 142] defined by the Floquet operator
@@
where $h_{j}$ is the on-site magnetic field, and $\ell=0,1$. For RD, system size drifts at the ETH-MBL crossover in KIM are milder than in autonomous spin chains [143]. To consider QP KIM, we set $h_{j}=\pi[1+\cos(2\pi\kappa j+\varphi)]$ with $\kappa=\kappa_{\mathrm{gr}}$ and select $\varphi\in(0,2\pi]$ with uniform probability. To tune the system across ETH-MBL crossover, we set $J=g=1/W$, with $W$ playing the role of the effective strength of QP potential. Our results for QP KIM will be contrasted with RD KIM, for which the onsite fields are selected with uniform probability from the interval $h_{j}\in[0,2\pi]$ and we still follow the choice $J=g=1/W$, with $W$ playing the role of disorder strength for RD KIM.
To show the generality of our findings, we consider also the Heisenberg spin-1/2 chain (HSC[62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74] by setting $\hat{H}_{0}=\frac{J}{4}\sum_{j=1}^{\ell}\vec{\sigma}_{j}\cdot\vec{\sigma}_{j+1}$ in Eq. (1) to arrive at RD HSC and in Eq. (2) to obtain the QP HSC. For QP systems (both Floquet and autonomous), due to the nature of QP potential, we consider open boundary conditions and set $\ell=1$. For RD we set periodic boundary conditions by putting $\ell=0$. We have verified that our conclusions apply also to RD systems with open boundary conditions.
Methods and observables.
@@
To study the ETH-MBL crossover in Floquet (autonomous) systems, we compute the eigenvectors $|\phi_{n}\rangle$ and the eigenvalues $e^{i\phi_{n}}$ ($E_{n}$) using the POLFED algorithm [46]. For the Floquet case, the algorithm relies on geometric sum spectral transformation [144, 143], allowing us to study system sizes up to $L=20$, substantially beyond the scope of traditional exact-diagonalization schemes, limited to $L\leq 14$ for unitary matrices.
@@
We study the properties of two ergodicity breaking indicators commonly employed in numerical analysis of the ETH-MBL crossover [145, 46]: the mean gap ratio and the entanglement entropy of eigenstates. The mean gap ratio is defined as $\overline{r}=\langle\min(g_{i},g_{i+1})/\max(g_{i},g_{i+1})\rangle$, where $g_{i}=\phi_{i+1}-\phi_{i}$ is the gap between two consecutive eigenphases for Floquet systems (while $g_{i}=E_{i+1}-E_{i}$ for HSC), and delimited-⟨$\langle$..$\rangle$ … ⟩ represents the average over eigenstates and disorder realizations. To characterize eigenvectors, we compute their entanglement entropy by splitting the system in two subsystems of equal size and evaluating $\mathcal{S}=-\sum_{i=1}^{i_{M}}\lambda_{i}^{2}\log\lambda_{i}^{2}$, where $\lambda_{i}$ are the Schmidt coefficients of the eigenstate $|\phi_{n}\rangle$ [146]. The entanglement entropies are averaged over eigenstates and disorder realizations, resulting in $\langle\mathcal{S}\rangle$. Subsequently, we rescale the result by the average entanglement entropy $\mathcal{S}_{\mathrm{COE}}$ of the eigenstates of COE matrices, i.e., $\overline{s}=\langle\mathcal{S}\rangle/\mathcal{S}_{\mathrm{COE}}$ [147, 148].
@@
In the following analysis, we compute $\mathrm{N_{ev}}=\min(2^{L}/10,1000)$ eigenvectors with eigenphases closest to $\phi_{n}=0.0$ for Floquet systems. For systems up to $L=16$, we consider at least $10^{4}$ realizations of the QP (or RD) potentials for each value of $W$, while for $L=17-18$ and $20$, we take at least $10^{3}$ and $200$ realizations of the system, respectively. For HSC we consider energies $E_{n}$ closest to the middle of the spectrum $E_{\mathrm{tg}}=(E_{\min}+E_{\max})/2$ [149].
The ETH-MBL crossover.
@@
Since MBL in Floquet systems with QP potential has not been analyzed in the literature so far, we begin by investigating the QP KIM. With the growth of QP potential strength $W$, the mean gap ratio $\overline{r}$, shown in Fig. 1 (a), interpolates between the value $\overline{r}_{\mathrm{COE}}\approx 0.53$ characteristic for ergodic systems [147] and $\overline{r}_{\mathrm{PS}}\approx 0.386$, signifying the MBL regime [27]. In the same disorder range, the rescaled entanglement entropy $\overline{s}$ changes from volume-law ($\overline{s}\approx 1$) to area-law ($\overline{s}\approx 0$[149].
@@
At first glance, the behavior of $\overline{r}$ is similar to the RD case. However, the ETH-MBL crossover also possesses certain unexpected features that are characteristic of the QP potential. Firstly, $\overline{r}$ is not changing smoothly with $W$ at small ($L\lesssim 10$) system sizes but rather fluctuates around a smooth curve monotonously decreasing with $W$. The fluctuations can be reduced by introducing randomness to the incommensurability factor $\kappa$, which can be taken from a narrow distribution centered at $\kappa_{\mathrm{gr}}$, see [149]. Hence, we fix $\kappa=\kappa_{\mathrm{gr}}$ for $L\geq 12$, while for smaller system sizes we take $\kappa$ uniformly distributed in the interval $[0.98,1.02]\times\kappa_{\mathrm{gr}}$. The second feature of ETH-MBL crossover characteristic for QP potential is that the crossover significantly sharpens up with increasing $L$.
Figure 1: The ETH-MBL crossover in QP KIM. (a) Mean gap ratio $\overline{r}$ as function of QP potential amplitude $W$ for various system sizes $L$, the dashed lines correspond to ergodic $\overline{r}_{\mathrm{COE}}$ and MBL $\overline{r}_{\mathrm{PS}}\approx 0.386$ limits; (b) Derivative $\alpha=d{\overline{r}}/dW$ as function of $W$, the red circles show the biggest slope $|\alpha_{\max}|$ at fixed $L$; Inset: exponential growth of $|\alpha_{\max}|$ with $L$.
To quantitatively analyze the sharpening up of the ETH-MBL crossover, we fit the mean gap ratio $\overline{r}$ in the interval $\overline{r}\in[\overline{r}_{\mathrm{COE}}-\delta_{\bar{r}},\overline{r}_{\mathrm{PS}}+\delta_{\bar{r}}]$ with a degree $5$ polynomial function of $W$ and evaluate its derivative $\alpha=d{\overline{r}}/dW$. Fixing $\delta_{\bar{r}}=0.01$, we obtain, for QP KIM, results shown in Fig. 1 (b). The derivative $\alpha$ admits a well-defined minimum at $W=W^{\alpha}_{\overline{r}}$. Notably, the corresponding absolute value of the slope $|\alpha_{\max}|$ is increasing, to good accuracy, exponentially with the system size, $|\alpha_{\max}|\propto e^{\lambda L}$ with $\lambda=0.093(4)$, see the inset in Fig. 1 (b). The robustness of this exponential scaling can be tested by looking at different intervals of $W$, polynomial functions, or fixed points of $W$, with all these methods giving a similar exponent $\lambda\approx 0.1$. The exponential sharpening up of the ETH MBL crossover, observed so far for QP KIM, constitutes the main result of this work. Before we further discuss this behavior in a broader context, we analyze the system size drifts at the ETH-MBL crossover in QP KIM.
System size drifts at the crossover.

### v2

# Many-body localization crossover is sharper in quasiperiodic potential

@@ Abstract
The strong disorder may significantly slow down or even completely hinder the thermalization of quantum many-body systems due to many-body localization (MBL). A sufficiently deep quasiperiodic potential may also inhibit thermalization. In this work, we numerically demonstrate direct differences in the behavior of standard ergodicity-breaking indicators at the MBL crossover in random and quasiperiodic systems. Our key finding is the exponential increase in the sharpness of the MBL crossover with system size for quasiperiodic systems, a trend that is only linear in disordered systems. The strong tendency towards a non-analytic behavior in quasiperiodic systems is consistent with the existence of dynamical regimes with sharply defined boundaries or an MBL phase transition. It highlights the importance of quasiperiodic systems for our understanding of many-body dynamics.
I Introduction
@@
where $\kappa$ is an incommensurability factor (e.g. $\kappa=\kappa_{\mathrm{gr}}\equiv(\sqrt{5}-1)/2$) and $\varphi$ is a global phase randomly selected from the interval $[0,2\pi]$ for each realization of the QP potential. Increase in the amplitude $W$ of the RD and QP potential leads to inhibition of thermalization and the onset of MBL regime [28, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92]. The QP potential is realized naturally in optical lattices by imposing a secondary standing wave with a period incommensurate with the primary lattice [93, 94, 95, 96, 97]. However, the RD and QP potentials are starkly different. The fields $h_{i}$ for RD are uncorrelated, while the QP potential (2) is deterministically fixed by the phase $\varphi$ for each realization, resulting in non-vanishing correlations between the on-site potential values at arbitrarily distant sites. Consequently, RD systems are more heterogeneous and host Griffiths type [98] rare regions of locally weaker disorder [99, 100, 101, 102, 103, 104, 105, 106, 107], while QP systems are more uniform [108].
@@
Do the disparities of the RD and QP potential translate to any essential differences in the properties of MBL in the two types of systems? The gross features of the dynamics in the MBL regime of QP systems [93, 94, 95, 96, 97] and of RD systems [109, 110, 111, 112] are similar. This includes a slow decay of local correlation functions [15, 94, 82, 52], a logarithmic in time growth of entanglement entropy both in RD  [20, 21, 22] and QP [113, 95] systems, and a long-lived memory of the initial state [114, 31, 84]. Moreover, the standard ergodicity breaking indicators such as the average gap ratio probing energy level statistics [27, 115] or rescaled entanglement entropy of eigenstates show qualitatively similar behavior at the ETH-MBL crossover both in RD and QP systems [116, 117]. Differences between RD and QP systems, suggested by a renormalization group schemes [118, 119], become apparent when quantities measuring sample-to-sample fluctuations of entanglement entropy [116], level statistics [120], and properties of self-energy term in the Fock space propagator [121] are considered. Such quantities indicate weaker fluctuations of the system’s properties for QP potential than for RD. Similarly, finite-size collapses point out weaker system size drifts for QP potential [117], simulations of time dynamics reveal weak but noticeable long-time oscillations in correlation functions of local observables for QP potential, unobservable for the RD case [52], while computations of diffusion constant indicate that it vanishes faster with growth of QP potential’s amplitude than in RD systems [122].
@@
In this work, we ask whether the differences between the QP and RD potentials imply differences in the behavior of the standard ergodicity-breaking indicators, such as the average gap ratio or probes of entanglement of eigenstates at the ETH-MBL crossover. To this end, we perform a large-scale numerical investigation of the ETH-MBL crossover in QP and RD Floquet systems. The latter are contrasted with results for RD, Eq. (1), and QP, Eq. (2), autonomous spin chains. Analyzing QP models, we find persistent drifts of ergodicity-breaking indicators at the ETH-MBL crossover, akin to the RD case. Despite these similarities, there is an immediate disparity between the QP and RD potentials. The ETH-MBL crossover sharpens up exponentially quickly with increasing system size $L$ for QP potential, while the sharpening up is only linear with $L$ for the RD. This demonstrates a direct difference between the standard bulk ergodicity-breaking indicators at ETH-MBL crossover in the RD and QP systems and constitutes the main finding of this work.
II Models and Numerical Methods
@@
Similarly to the autonomous spin chains, periodically driven Floquet systems may undergo MBL [123, 124, 125, 126, 127, 128, 129, 130, 131] which prevents their heating [132], enabling creation of driven time crystals [133, 134, 135, 136, 137, 138, 139] and stabilization of Floquet insulators [140, 141, 142, 143, 144]. We focus on a Kicked Ising Model (KIM[145, 146] defined by the Floquet operator
@@
where $h_{j}$ is the on-site magnetic field and $\ell$ determines the boundary conditions of the system. For QP systems due to the nature of QP potential, we consider open boundary conditions and set $\ell=L-1$. For RD we set periodic boundary conditions by putting $\ell=L$ and identifying $j=L+1$ with $j=1$. We have verified that our conclusions apply also to RD systems with open boundary conditions. For RD, system size drifts at the ETH-MBL crossover in KIM are milder than in autonomous spin chains [147]. To consider QP KIM, we set $h_{j}=\pi[1+\cos(2\pi\kappa j+\varphi)]$ with $\kappa=\kappa_{\mathrm{gr}}$ and select $\varphi\in(0,2\pi]$ with uniform probability. To tune the system across ETH-MBL crossover, we set $J=g=1/W$, with $W$ playing the role of the effective strength of QP potential. Our results for QP KIM will be contrasted with RD KIM, for which the onsite fields are selected with uniform probability from the interval $h_{j}\in[0,2\pi]$ and we still follow the choice $J=g=1/W$, with $W$ playing the role of disorder strength for RD KIM.
To show the generality of our findings, we consider also the Heisenberg spin-1/2 chain (HSC[62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74] by setting
$\hat{H}_{0}=\frac{J}{4}\sum_{j=1}^{\ell}\vec{\sigma}_{j}\cdot\vec{\sigma}_{j+1}$ (4)
in Eq. (1) to arrive at RD HSC and in Eq. (2) to obtain the QP HSC.
@@
To study the ETH-MBL crossover in Floquet (autonomous) systems, we compute the eigenvectors $|\phi_{n}\rangle$ and the eigenvalues $e^{i\phi_{n}}$ ($E_{n}$) using the POLFED algorithm [46]. For the Floquet case, the algorithm relies on geometric sum spectral transformation [148, 147], allowing us to study system sizes up to $L=20$, substantially beyond the scope of traditional full exact-diagonalization schemes, limited, on present-day computers, to $L\leq 14$ for unitary matrices.
We study the properties of two ergodicity-breaking indicators commonly employed in the numerical analysis of the ETH-MBL crossover [149, 46]: the mean gap ratio and the entanglement entropy of eigenstates. The mean gap ratio is defined [27] as
$\overline{r}=\langle\min(g_{i},g_{i+1})/\max(g_{i},g_{i+1})\rangle$ (5)
@@
where $g_{i}=\phi_{i+1}-\phi_{i}$ is the gap between two consecutive eigenphases for Floquet systems (while $g_{i}=E_{i+1}-E_{i}$ for HSC), and delimited-⟨$\langle$..$\rangle$ … ⟩ represents the average over eigenstates and disorder realizations. To characterize eigenvectors, we compute their entanglement entropy by splitting the system in two subsystems and evaluating
$\mathcal{S}=-\sum_{i=1}^{i_{M}}\lambda_{i}^{2}\log\lambda_{i}^{2}$ (6)
where $\lambda_{i}$ are the Schmidt coefficients of the eigenstate $|\phi_{n}\rangle$ [150]. The entanglement entropies are averaged over eigenstates and disorder realizations, resulting in $\langle\mathcal{S}\rangle$. Subsequently, we rescale the result by the average entanglement entropy $\mathcal{S}_{\mathrm{COE}}$ of the eigenstates of Circular Orthogonal Ensembles (COE) matrices [151], i.e., $\overline{s}=\langle\mathcal{S}\rangle/\mathcal{S}_{\mathrm{COE}}$ [152, 153].
As an additional measure of entanglement of eigenstates we use the mean Schmidt gap, $\overline{\Delta}$, obtained by averaging Schmidt gaps $\Delta=\lambda_{1}^{2}-\lambda_{2}^{2}$ of eigenstates. One expects the average Schmidt gap to be decaying exponentially with system size $L$ for the eigenstates in the ETH regime, while in the MBL regime, with low-entanglement entropy, the mean Schmidt gap is expected to be independent of system size [154].
@@
In the following analysis, we compute $\mathrm{N_{ev}}=\min(2^{L}/10,1000)$ eigenvectors with eigenphases closest to $\phi_{n}=0.0$ for Floquet systems. For systems up to $L=16$, we consider at least $10^{4}$ realizations of the QP (or RD) potentials for each value of $W$, while for $L=17-18$ and $20$, we take at least $10^{3}$ and $200$ realizations of the system, respectively.
Figure 1: Top panel: The ETH-MBL crossover in QP KIM showed by typical ergodicity breaking indicators as a function of the effective quasi-periodic (QP) potential strength $W$ for various system sizes $L$. These indicators include the (a) average gap ratio $\overline{r}$, (b) the rescaled entanglement entropy $\overline{s}$, and (c) the Schmidt gap $\overline{\Delta}$. Lower panel: Their respective derivatives $\alpha=d{\overline{x}}/dW$ as function of $W$, where $\overline{x}$ is the respective observable. The red circles highlighting the maximum slope $|\alpha_{\max}|$ for each $L$. Inset: A semi-logarithmic plot of $|\alpha_{\max}|$ versus $L$, illustrating an exponential scaling behavior.
III The ETH-MBL crossover in QP Kicked Ising Model
@@
While Floquet systems with RD were quite extensively studied, as reviewed above, it is not so with QP potential case. Apart of the early experiment on Floquet MBL in QP potential [155] the theoretical studies concentrated mostly on multifractality of eigenstates [156, 86]. We shall consider the KIM in QP potential having the data for RD potential at hand [147]. With the growth of QP potential strength $W$, the mean gap ratio $\overline{r}$, shown in Fig. 1 (a), interpolates between the value $\overline{r}_{\mathrm{COE}}\approx 0.53$ characteristic for ergodic systems [152] and $\overline{r}_{\mathrm{PS}}\approx 0.386$, signifying the MBL regime [27]. In the same disorder range, the rescaled entanglement entropy $\overline{s}$ changes from volume-law ($\overline{s}\approx 1$) to area-law ($\overline{s}\approx 0$) as shown in Fig. 1 (b). In the same spirit, the ETH-MBL crossover is also captured by the Schmidt gap $\overline{\Delta}$ which distinguishes the ergodic regime in which $\overline{\Delta}$ decreases with increase of $L$ from the MBL regime at large $W$, where $\overline{\Delta}$ is approximately independent of $L$, as shown in Fig. 1 (c).
@@
At first glance, the behavior of these standard ergodicity breaking indicators is similar to the RD case [27, 28, 30, 154, 46]. However, the analyzed ETH-MBL crossover in QP KIM also possesses certain unexpected features characteristic for the QP systems. Firstly, $\overline{r}$ is not changing smoothly with $W$ at small ($L\lesssim 10$) system sizes but rather fluctuates around a smooth curve monotonously decreasing with $W$. The fluctuations can be reduced by introducing randomness to the incommensurability factor $\kappa$, which can be taken from a narrow distribution centered at $\kappa_{\mathrm{gr}}$, see Appendix A. Hence, we fix $\kappa=\kappa_{\mathrm{gr}}$ for $L\geq 12$, while for smaller system sizes we take $\kappa$ uniformly distributed in the interval $[0.98,1.02]\times\kappa_{\mathrm{gr}}$. The second feature of ETH-MBL crossover characteristic for QP potential is that the crossover significantly sharpens up with increasing $L$.
To quantitatively analyze the sharpening up of the ETH-MBL crossover, we fit the observable $\overline{x}$ (with $\overline{x}\in\{\overline{r},\overline{s},\overline{\Delta}\}$) in the interval ${\overline{x}}\in[{\overline{x}_{\mathrm{COE}}}-\delta_{{\overline{x}}},{\overline{x}}_{\mathrm{PS}}+\delta_{{\overline{x}}}]$ with a degree $5$ polynomial function of $W$ and evaluate its derivative $\alpha=d{{\overline{x}}}/dW$. The choice $\delta_{\overline{x}}$ depends on which observable we are looking at. For the average gap ratio, we take $\delta_{\overline{r}}=0.01$, while for the rescaled entanglement entropy and the Schmidt gap $\delta_{\overline{x}}=0.1$. We have checked that the results are not too sensitive to the values of $\delta_{\overline{x}}$ chosen, see Appendix B. Interestingly, the corresponding absolute value of the slope $|\alpha_{\max}|$, we obtain, is increasing, to good accuracy, exponentially with the system size, as shown in the insets of the lower panels in Fig. 1 (d, e, f). Parametrizing the exponential dependence as $|\alpha_{\max}|=Ae^{\lambda L}$, where $A,\lambda$ are constants, we obtain, for the average gap ratio, $\overline{r}$, $\lambda=0.093(4)$, while for the rescaled entanglement entropy, $\overline{s}$ we get $\lambda=0.068(2)$ and for the Schmidt gap, $\overline{\Delta}$, we find $\lambda=0.064(2)$.
The robustness of the exponential scaling of $|\alpha_{\max}|$ can be tested by looking at different intervals of $W$, polynomial functions, or fixed points of $W$, with all these methods giving a similar exponent $\lambda\approx 0.1$, as we further discuss in the Appendix B. While the uncovered increase of $|\alpha_{\max}|$ is clearly faster than linear in $L$, the dependence of $|\alpha_{\max}|$ on $L$ could also be fitted by a second, or higher order polynomial of $L$. However, such fits would involve at least three fitting parameters. In contrast, the exponential system size dependence, $|\alpha_{\max}|=Ae^{\lambda L}$, is the simplest, two-parameter formula that accurately describes the behavior of $|\alpha_{\max}|$ for QP KIM in the interval of system sizes accessible to us. With this remark in mind, we keep referring to the behavior of the $|\alpha_{\max}|$ in QP KIM as the exponential growth.
The behavior of the parameter $\lambda$ is robust against different choices of the disorder strength at which the value of $d\overline{x}/dW$ is analyzed. For instance, instead of considering the disorder strength corresponding to the maximum of the absolute value of the derivative, let us consider the slopes, $\alpha$, at the disorder strength at which the ergodicity-breaking indicators admit a value $\overline{x}_{\mathrm{avg}}$ in the middle between the $\overline{x}$ values for the ETH and MBL limits. For the average gap ratio $\overline{r}$, the slopes at these points are denoted by blue cross markers in Fig. 2 (a). Initially, the slope at this point is far from $\alpha_{\mathrm{max}}$ (red circles), but as the system size is increased, the slopes approach each other. Both choices exhibit the exponential scaling of $\alpha$ with the system size, as illustrated in Fig. 2 (b), featuring a similar exponent $\lambda\approx 0.1$. Similar conclusions are drawn for the rescaled entanglement entropy $\overline{s}$ and the Schmidt gap $\Delta$ (data not shown).
This consistency of the behavior of $\lambda$ for different choices of the point at which the slope is evaluated confirms the robustness of the exponential scaling of $|\alpha|$ at the ETH-MBL crossover in QP KIM, which is the main result of this work. However, before we further discuss this behavior in a broader context, we analyze the system size drifts at the ETH-MBL crossover in QP KIM.
Figure 2: (a) Derivatives of $\overline{r}(W)$ for different system sizes across the MBL crossover, where the red circles represent the biggest slope and the blue cross is the slope value at $\overline{r}_{\mathrm{avg}}\approx 0.458$; (b) A direct comparison between the exponential behavior of these two different points. Despite minor differences, both curves scale with similar exponent ($\lambda\approx 0.1$).
III.1 System size drifts at the ETH-MBL crossover
To investigate the finite-size drifts at the ETH-MBL crossover and shed light on possible behavior of the crossover in the large $L$ limit, we follow [46] and consider two different size-dependent disorder strengths:
  • $W_{\overline{x}}^{T}(L)$: the disorder strength at which the ergodicity breaking indicator $\overline{x}(W)$, at a given $L$, deviates from its ergodic value by a small number $\delta_{\bar{x}}$;
  • $W_{\overline{x}}^{*}(L)$: the crossing point between the curves $\bar{x}(W)$ for two different system sizes, $L-\Delta L$ and $L+\Delta L$, where $\Delta L\ll L$.